# Their currents turn awry, and lose the name of action. I: Fundamental limits to orbit reconstruction due to non-conservation of stellar actions

### v1

@@ Abstract
pubyear: 2025pagerange: Their currents turn awry, and lose the name of action. I: Fundamental limits to orbit reconstruction due to non-conservation of stellar actionsA
@@ 1 Introduction
Most existing studies of action evolution over time focus on the global distribution of actions, particularly in the context of radial migration and dynamical heating. Simulations have explored how secular evolution redistributes angular momentum and heats the disc over gigayear timescales (Roškar et al., 2008; Vera-Ciro et al., 2014; Monari et al., 2016; Vera-Ciro & D’Onghia, 2016; Halle et al., 2018; Mikkola et al., 2020; Okalidis et al., 2022). Observational studies, on the other hand, study action evolution by comparing the present-day properties of stars of different ages, often in the context of, again, disc heating and radial migration (Frankel et al., 2018; Ting & Rix, 2019; Frankel et al., 2020). However, these studies primarily describe population-wide trends, while what matters for reconstruction studies is the time evolution of the individual stars’ actions. This is analogous to the distinction between the microscopic motion of gas molecules and the macroscopic evolution of temperature in a thermodynamic system – such a system may have a completely constant, predictable temperature, yet it may still be impossible to backtrace the trajectory of any individual molecule over any substantial period.
@@ 3.1 Gravitational potential profile
To accurately model the time-dependent gravitational potential of the galaxy, we first define an appropriate reference frame. We cannot simply use the simulation frame because, although the simulation is initialised with the galactic plane at rest at $z=0$ and cylindrically symmetric about the origin, over the duration of the simulation supernova explosions drive an asymmetric wind off the galaxy, imparting a non-negligible momentum to the remaining gas. To correct for this shift, and thereby ensure that we are calculating our actions with respect to the rest frame of the galaxy, we calculate the centre of mass (COM) of cold gas particles (those with temperature $\leq 10^{4}$ K) at each snapshot; we plot the resulting position in $z$-direction as a function of time in Figure 2. We observe a velocity $\sim 2$ km/s in the $z$ direction and $\sim 1$ km/s in the $x$ and $y$ directions, and this movement of the COM of the cold gas is consistent with what we expect: the simulation produces galactic winds with a mass flux $\sim 1$ $M_{\odot}$/yr and a velocity of a few hundred km/s (WK23), so over the simulation timescale of $\sim 500$ Myr winds eject $\sim 5\times 10^{8}\mbox{$M_{\odot}$}$, roughly 10% of the total cold gas mass $\approx 4.7\times 10^{9}\mbox{$M_{\odot}$}$ at the start of the simulation. If the winds were completely one-sided, this would therefore be sufficient to accelerate the gas disc to several tens of km s-1, but since the winds are only somewhat asymmetric some of the momentum cancels. In addition, gravitational forces transfer some of the momentum from the gas to the stellar disc, which is roughly ten times as massive. Consequently, the net velocity is reduced to only a few km/s. The $z$ velocity is greater than the $x$ and $y$ velocities due to the greater escape of winds normal to the galactic plane.
@@
Figure 2: Coordinates of the centre of mass of cold gas in the simulation with time in the $z$ direction. We also show a least-squares linear fit to the $z$ position (solid line), the functional form for which is provided in the legend (with position in units of pc and times in units of Myr); the $R^{2}$ value for the fit is 0.94.
Having established our reference frame, we are now ready to compute our best estimate of an azimuthally symmetric galactic potential from large but sparse samples at our star particle positions. To do so, for each simulation snapshot we define a grid in cylindrical coordinates that spans the entire galaxy, with radial coordinates extending from 0.1 pc to 20 kpc, a vertical range of ±$\pm 1$ kpc, and a resolution of 1 pc in both directions. The grid in the azimuthal direction has a resolution of $252\pi/$ 25 radians. For each point on the defined grid, we identify the nearest-neighbour particle and assign its potential value (which is computed by the gizmo gravity solver) to that grid point. We then average over the $\phi$ direction to create an axisymmetric 2D potential grid in $(R,z)$. Figure 3 shows some sample slices through the 2D potential we derive for one time snapshot.
@@ 3.2 Action calculation
Our next step is to calculate the actions of each star. Performing this calculation in full generality requires computationally-expensive numerical integration. However, we are solely concerned with stars that are both young – age $\lesssim 0.5$ Gyr – and near the galactic plane – root mean square height $\langle z^{2}\rangle^{1/2}<200$ pc, and $|z|<1$ kpc for all stars. For stars with these properties, the epicyclic approximation, whereby we decompose stellar orbits into independent radial and vertical oscillations about a guiding center, is highly accurate (Hunt & Vasiliev, 2025). Quantitatively, Solway et al. (2012) find that, even in the presence of spiral structures, for stars up to 1 kpc off the plane, and over a timescale of a few Gyr, vertical actions computed using the epicyclic approximation change by 20.7%, compared to 15.6% with a more exact calculation; this small difference confirms that the epicyclic approximation is appropriate for our purposes. This is advantageous, because in the epicyclic approximation actions are far cheaper to compute.
@@
We compute $R_{g}$, $\kappa$ and $\nu$ for each star as follows. We first use the star’s radial position $R$ and azimuthal velocity $v_{\phi}$ to evaluate its angular momentum $L_{z}=Rv_{\phi}$, and then to use Equation 4 together with the galactic rotation curve $\Omega(R)$ to evaluate the star’s guiding radius $R_{g}$. Once $R_{g}$ is known, we can then evaluate $\kappa(R_{g})$ and $\nu(R_{g})$ from Equation 6 and Equation 7 using the known rotation curve and potential. The primary challenge to executing this strategy is that the gravitational potential from the simulation is discrete and subject to numerical noise, and thus considerable care is required when evaluating the derivatives that appear in the expressions above. We provide a detailed description of our full methodology in Appendix A.
@@
Figure 4: An example of a star that exhibits rapid variation in actions, likely because it is part of a star cluster. The top two panels show the $R$ and $z$ position as a function of time, while the bottom three show action versus time. The grey lines represent the actual actions we compute, while the pink lines show the actions after applying the low-pass filter as discussed in subsection 3.3.
3.3 Action evolution study
Once we have computed the actions for all stars at each snapshot, we are in a position to investigate their temporal evolution. Initial analysis reveals that a significant fraction of stars exhibit rapid variations in their actions over short (few Myr or less) timescales, likely the result of being part of gravitationally bound structures such as star clusters; we discuss these structures further in section 5, and they will be the principal focus of Paper II in this series. For the moment we wish to focus on the longer term secular evolution of actions, and to isolate these from the high-frequency variations we apply a Butterworth filter with a cut-off frequency of 1/30 Myr-1 to the time series of actions, effectively suppressing variations occurring on timescales shorter than 30 Myr. Figure 4 illustrates an example of a star displaying rapid variations in its actions, along with the impact of the low-pass filter in smoothing these fluctuations.
@@ 3.3 Action evolution study
where $i=(R,z,\phi)$ is the component, $t$ is the time of the snapshot, $\Delta t$ is the time interval between the two snapshots and $J_{i}(t)$ is the $i$ component of the star’s action action at time $t$ in the simulation. Similarly, we define the relative difference as
$\delta_{i}(\Delta t)=\frac{\Delta J_{i}(t,\Delta t)}{J_{i}(t)}.$ (16)
@@
We have $\approx 450$ time snapshots, and therefore $\approx 10^{5}$ snapshot pairs, and $\sim 10^{6}$ stars, so our total sample for analysis consists of $\sim 10^{11}$ action differences.
4 Results
We are now prepared to characterise the time-evolution of the actions. We first examine the evolution of the distribution of actions in subsection 4.1, and then study the properties of the time series describing individual stars rather than the full population in subsection 4.2. We discuss how the evolution varies with the environmental conditions under which stars are born in subsection 4.3 and subsection 4.4. In subsection 4.5, we compare the action evolution for two different population of stars.
4.1 Distribution of actions
Before examining the evolution of individual stellar actions over time, we first study the distribution of stellar actions and its time evolution. We do so by selecting our final snapshot ($t=464$ Myr) and examining the action distribution as a function of stellar age, since this provides a snapshot of the data that can be compared directly to observations. We place all stars at the last snapshot in 1 kpc-wide bins of radial position $R$, and in Figure 5 we show the distributions of radial ($J_{R}$), azimuthal ($J_{\phi}$), and vertical ($J_{z}$) actions for stars as a function of stellar age for the 4-5 kpc, 9-10 kpc, and 12-13 kpc radial bins. The solid central lines represent the medians of these distributions, while the shaded areas show the 16th to 84th percentile range.
Our results reveal that the overall distribution of actions remains relatively stable over the simulation period. We find that $J_{\phi}$ is essentially static, which is not surprising, given that our $\approx 500$ Myr run time is too short for significant radial mixing. There is some weak evolution of $J_{R}$ and $J_{z}$ distributions, and the latter is particularly useful because there have been several observational studies of this evolution to which we can compare. Ting & Rix (2019) examine this quantity for stars of age up to 8 Gyr and observe significant broadening of the distribution, but this timescale is much larger than that for which our simulations run and is therefore not directly comparable. More recently, Garzon et al. (2024) performed a similar analysis for stars younger than 0.5 Gyr, well-matched to the timescale of our simulation. We overlay their results in the bottom row of our plot (in red). Since their analysis is limited to stars with the Galactocentric radii within 8 to 13 kpc, we only have comparisons for two of the radial bins plotted. For the 9-10 kpc radial bin, we find that our results agree very well with their mean $J_{z}$ values across ages. However, in the 12-13 kpc radial bin, we find significantly lower actions and less evolution. Garzon et al. (2024) suggest that the very large change in action evolution from $9-10$ kpc to $12-13$ kpc that they find may be due to the warping of the outer Milky Way disc (e.g., Uppal et al., 2024, and references therein), a feature that is absent from our simulations, since we begin with a flat disc and do not include perturbations from dwarf galaxies or dark matter sub-halos that might induce a warp. This is likely the reason that we do not reproduce this aspect of the observations in our simulations. This also means that the results we derive below for the rates at which individual stellar actions evolve are likely to be underestimates in the far outer disc beyond $\sim 10$ kpc where the warp is significant. We should therefore think of the results we obtain as lower limits that describe action evolution in the absence of large-scale external perturbations to the galactic potential.
4.2 Change in actions over time
@@
Figure 6: Heatmap of the median of the absolute change in stellar actions (row-wise in order: $\Delta J_{R}$, $\Delta J_{z}$, and $\Delta J_{\phi}$) as a function of the time lag and stellar age. The dotted lines indicate stellar ages of 5, 10, 50, 100, 200, 300 and 400 Myr, corresponding to the time slices plotted in Figure 7.
@@
We now investigate the evolution of individual stellar actions over time as characterised by the absolute changes $\Delta J_{R}$, $\Delta J_{z}$ and $\Delta J_{\phi}$ (Equation 15). These quantities are functions of both the current stellar age $t_{*}$ and the time lag $\Delta t$, and we therefore place all stars in 2D bins of these quantities and compute the median in each bin; we use medians rather than means to ensure robustness against outliers. We show the result in Figure 6; in this plot, stellar age appears on the $x-$axis and the time lag $\Delta t$ on the $y-$axis. The expected behaviour is immediately visible – small variations at short time intervals appear as the darkest colours in a horizontal band at the bottom of the plot. We see only a weak dependence on stellar age, as there are no very obvious vertical features in the heatmap.
@@
To take a closer look, we extract vertical slices from the heatmap at ages 5, 10, 50, 100, 200, 300 and 400 Myr, as shown by the dashed lines in Figure 6. These slices, illustrated in Figure 7, show how the absolute change in actions depends on the time lag for stars of different ages. We see here that there is a weak dependence of $\Delta J$ on stellar age, but in a direction that is opposite what one might naively expect – we expect younger stars to exhibit larger changes in their actions than older stars because they are more likely to reside near their natal molecular clouds (Lada & Lada, 2003), where ongoing interactions with dense gas or other stars can lead to greater dynamical perturbations. Conversely, older stars have typically migrated away from their birth regions and are less influenced by such perturbations. Contrary to these expectations, we instead find that older stars undergo more rapid changes to their actions, as indicated by the larger slopes in Figure 7.
@@
Figure 7: Median of the absolute change in stellar actions (row-wise in order: $\Delta J_{R}$, $\Delta J_{z}$, and $\Delta J_{\phi}$) for stars of different ages across the simulation as a function of time lag. Different stellar age groups are shown in different colours as indicated in the legend.
@@
To understand why this occurs, we next examine relative rather than absolute changes in actions, focusing on $J_{R}$ and $J_{z}$ since $J_{\phi}$ is dominated by circular orbital motion and is less sensitive to dynamical perturbations. Figure 8 shows the median square relative changes of these two actions, $\delta J_{R}^{2}$ and $\delta J_{z}^{2}$, as a function of time lag for the same stellar ages as shown in Figure 7. In this plot, we see that the different stellar age bins collapse onto nearly a single line, suggesting a universal behaviour among stars of all ages.

### v2

# Fundamental limits to orbit reconstruction due to non-conservation of stellar actions in a Milky Way-like simulation

@@ Abstract
pubyear: 2025pagerange: Fundamental limits to orbit reconstruction due to non-conservation of stellar actions in a Milky Way-like simulationB
@@ 1 Introduction
Most existing studies of action evolution over time focus on the global distribution of actions, particularly in the context of radial migration and dynamical heating. Simulations have explored how secular evolution redistributes angular momentum and heats the disc over gigayear timescales (Roškar et al., 2008; Vera-Ciro et al., 2014; Monari et al., 2016; Vera-Ciro & D’Onghia, 2016; Halle et al., 2018; Mikkola et al., 2020; Okalidis et al., 2022). Observational studies similarly focus on population-wide properties, most notably through the age-velocity dispersion relation (AVR, e.g. Sharma et al., 2014; Yu & Liu, 2018; Mackereth et al., 2019), which is interpreted as evidence of cumulative dynamical heating from spiral structure, GMCs and other perturbations. A few observational studies have also studied action evolution by comparing the present-day actions of stars of different ages, again primarily in the context of disc heating and radial migration (Frankel et al., 2018; Ting & Rix, 2019; Frankel et al., 2020). However, these studies mainly characterise ensemble behaviour – how the statistical distribution of actions or velocities broadens with time – rather than the dynamical evolution of individual stellar actions. For applications such as orbit reconstruction, it is this individual time evolution of actions that is crucial. This is analogous to the distinction between the microscopic motion of gas molecules and the macroscopic evolution of temperature in a thermodynamic system – such a system may have a completely constant, predictable temperature, yet it may still be impossible to backtrace the trajectory of any individual molecule over any substantial period.
@@ 3.1 Gravitational potential profile
To accurately model the time-dependent gravitational potential of the galaxy, we first define an appropriate reference frame. We cannot simply use the simulation frame because, although the simulation is initialised with the galactic plane at rest at $z=0$ and cylindrically symmetric about the origin, over the duration of the simulation supernova explosions drive an asymmetric wind off the galaxy, imparting a non-negligible momentum to the remaining gas and to the stars to which it is gravitationally bound.
To correct for this shift, we calculate our actions with respect to the rest frame of the galaxy. To this end, we use the population of stellar disc particles that are already present in the initial conditions of Z25 (as described in Section 2.2). We refer to these pre-existing particles as ‘initial stars’. Having calculated the centre of mass (COM) of the initial stars at each snapshot, we plot the resulting position in $z$-direction as a function of time in Figure 2.
We observe a velocity $\sim 2$ km/s in the $z$ direction and $\sim 1$ km/s in the $x$ and $y$ directions, and this movement of the COM of the initial stars is consistent with what we expect: the simulation produces galactic winds with a mass flux $\sim 1$ $M_{\odot}$/yr and a velocity of a few hundred km/s (WK23), so over the simulation timescale of $\sim 500$ Myr winds eject $\sim 5\times 10^{8}\mbox{$M_{\odot}$}$, roughly 2% of the total initial stars’ mass $\approx 3\times 10^{10}\mbox{$M_{\odot}$}$. This results in a net velocity of the stellar and gas disc of order of a few km/s. The $z$ velocity is greater than the $x$ and $y$ velocities due to the greater escape of winds normal to the galactic plane.
@@
Figure 2: Coordinates of the centre of mass of initial stars in the simulation with time in the $z$ direction. We also show a least-squares linear fit to the $z$ position (solid line), the functional form for which is provided in the legend (with position in units of pc and times in units of Myr); the $R^{2}$ value for the fit is 0.94.
Having established our reference frame, we are now ready to compute our best estimate of an azimuthally symmetric galactic potential from large but sparse samples at our star particle positions. To do so, for each simulation snapshot we define a grid in cylindrical coordinates that spans the entire galaxy, with radial coordinates extending from 0.1 pc to 20 kpc, a vertical range of $\pm 1$ kpc, and a resolution of 1 pc in both directions. The grid in the azimuthal direction has a resolution of $2\pi/25$ radians. For each point on the defined grid, we identify the nearest-neighbour particle and assign its potential value (which is computed by the gizmo gravity solver) to that grid point. We then average over the $\phi$ direction to create an axisymmetric 2D potential grid in $(R,z)$. Figure 3 shows some sample slices through the 2D potential we derive for one time snapshot.
@@ 3.2 Action calculation
Our next step is to calculate the actions of each star. Performing this calculation in full generality requires computationally-expensive numerical integration, an approach that is impractical given the size of our data set, $\sim 10^{8}$ stellar particles. However, we are solely concerned with stars that are both young – age $\lesssim 0.5$ Gyr – and near the galactic plane – root mean square height $\langle z^{2}\rangle^{1/2}<200$ pc, and $|z|<1$ kpc for all stars. For stars with these properties, the epicyclic approximation, whereby we decompose stellar orbits into independent radial and vertical oscillations about a guiding center, is highly accurate (Hunt & Vasiliev, 2025). Quantitatively, Solway et al. (2012) find that, even in the presence of spiral structures, for stars up to 1 kpc off the plane, and over a timescale of a few Gyr, vertical actions computed using the epicyclic approximation change by 20.7%, compared to 15.6% with a more exact calculation; this small difference confirms that the epicyclic approximation is reliable for stars near the plane, particularly over the timescales of interest in our study. To further validate the suitability of this approximation for our dataset, in Appendix A we compare actions for a representative sample of stars computed using the epicyclic approximation to results we obtain using a more precise, but computationally intensive, orbit integration method. This comparison confirms that the errors introduced by the epicyclic approximation are small compared to the signals we measure.
Figure 4: An example of a star that exhibits rapid variations in actions, likely because it is part of a bound star cluster. The top three panels show the $R$ and $z$ position and $v_{\phi}$ velocity as a function of time, while the bottom three show actions versus time. The grey lines represent the actual actions we compute, while the pink lines show the actions after applying the low-pass filter as discussed in Section 3.2.
@@
We compute $R_{g}$, $\kappa$ and $\nu$ for each star as follows. We first use the star’s radial position $R$ and azimuthal velocity $v_{\phi}$ to evaluate its angular momentum $L_{z}=Rv_{\phi}$, and then to use equation 4 together with the galactic rotation curve $\Omega(R)$ to evaluate the star’s guiding radius $R_{g}$. Once $R_{g}$ is known, we can then evaluate $\kappa(R_{g})$ and $\nu(R_{g})$ from equation 6 and equation 7 using the known rotation curve and potential. The primary challenge to executing this strategy is that the gravitational potential from the simulation is discrete and subject to numerical noise, and thus considerable care is required when evaluating the derivatives that appear in the expressions above. We provide a detailed description of our full methodology in Appendix B.
@@
Figure 5: An example of a bound cluster in our simulation. We plot the $x$ and $y$ coordinates of the member stars (blue points) relative to the cluster’s centre of mass, showing their spatial configuration in six different snapshots. The snapshot time $t$ and the mean stellar age of the stars shown is annotated in each panel. Grey points show gas particles.
Once we have computed the actions for all stars at each snapshot, we are in a position to investigate their temporal evolution. Initial analysis reveals that a significant fraction of stars exhibit rapid variations in their actions over short (few Myr or less) timescales. These fluctuations are not caused by large changes in radial or vertical position but by substantial variations in azimuthal velocity $v_{\phi}$, as shown in Figure 4, which occur because the stars that exhibit this behaviour are part of gravitationally bound star clusters. We illustrate an example of such a bound structure in Figure 5. We identify cluster members as follows: we start with the example star that exhibits rapid action variations (in this case the one shown in Figure 4), and then we iteratively add the next-nearest star and check whether the resulting system is gravitationally bound by computing the total energy (kinetic + potential energy) of the system. We continue adding stars in increasing order of their distance from the original star until the inclusion of a new member renders the total energy positive. We do this at the last snapshot. We then track the resulting set of bound members backwards in time, and re-evaluate their boundedness at each of the snapshots shown in Figure 5. Thus the stars shown in each panel consist of those that are bound to our example star both at the time shown and at the final snapshot, and the persistence of large numbers of members between times indicates that we are examining a gravitationally bound population that remains together for many internal dynamical times. All of the stars we have examined that exhibit rapid action variations prove to be parts of structures such as this. In these gravitationally bound clusters, internal orbital motion about the cluster’s centre of mass drives coherent velocity changes in the galactocentric frame that we see in Figure 4. These internal dynamics, while physically real, are not indicative of long-term secular evolution in the galactic potential which is the primary focus of this study. Such bound structures will be the principal focus of Paper II in this series. For the moment, to isolate the long-term behaviour from the high-frequency variations, we apply a Butterworth filter with a cut-off frequency of 1/30 Myr-1 to the time series of actions, thereby suppressing variations occurring on timescales shorter than 30 Myr. As shown in Figure 4, this effectively smooths out the short-timescale oscillations caused by internal cluster dynamics while preserving secular trends in actions.
4 Simulation Validation
Before we examine how the actions of individual stars evolve – the primary purpose of this paper – we first validate our simulation and our method for calculating actions by comparing to past studies that have focused on the evolutions of actions or closely related quantities for stellar populations rather than individual stars. We begin in Section 4.1 by examining how the distribution of orbital actions for stars of different ages in our simulation compare to observations. We then perform a similar check in Section 4.2 by comparing the trend of stellar velocity dispersion (which is coarsely related to the action distribution) with age in our simulations to previous results.
4.1 Distribution of actions
In order to compare the action evolution in our simulations to observations, we select our final snapshot ($t=464$ Myr) and bin the stars in that snapshot by radial position $R$ and age $\tau$ (defined as $\tau=t-t_{\mathrm{form}}$, where $t_{\mathrm{form}}$ is the time at which each star formed); this provides a snapshot of the data that can be compared directly to observations. In Figure 6 we show the distributions of radial ($J_{R}$), azimuthal ($J_{\phi}$), and vertical ($J_{z}$) actions as a function of $\tau$ for the 4-5 kpc, 9-10 kpc, and 12-13 kpc radial bins. The solid central lines represent the medians of these distributions, while the shaded areas show the 16th to 84th percentile range.
Our results reveal that the overall distribution of actions remains relatively stable over the simulation period. We find that $J_{\phi}$ is essentially static, which is not surprising, given that our $\approx 500$ Myr run time is too short for significant radial mixing. There is some weak evolution of $J_{R}$ and $J_{z}$ distributions, and the latter is particularly useful because there have been several observational studies of this evolution to which we can compare. Ting & Rix (2019) examine this quantity for stars of age up to 8 Gyr and observe significant broadening of the distribution, but this timescale is much larger than that for which our simulations run and is therefore not directly comparable. More recently, Garzon et al. (2024) performed a similar analysis for stars younger than 0.5 Gyr, well-matched to the timescale of our simulation. We overlay their results in the bottom row of our plot (in red). Since their analysis is limited to stars with Galactocentric radii from 8 to 13 kpc, we only have comparisons for two of the radial bins plotted. For the 9-10 kpc radial bin, we find that our results agree very well with their mean $J_{z}$ values across ages. However, in the 12-13 kpc radial bin, we find significantly lower actions and less evolution. Garzon et al. (2024) suggest that the very large change in action evolution from 9109-10 kpc to 121312-13 kpc that they find may be due to the warping of the outer Milky Way disc (e.g., Uppal et al., 2024, and references therein), a feature that is absent from our simulations, since we begin with a flat disc and do not include perturbations from dwarf galaxies or dark matter sub-halos that might induce a warp. This is likely the reason that we do not reproduce this aspect of the observations in our simulations. This also means that the results we derive below for the rates at which individual stellar actions evolve are likely to be underestimates in the far outer disc beyond $\sim 10$ kpc where the warp is significant. We should therefore think of the results we obtain as lower limits that describe action evolution in the absence of large-scale external perturbations to the galactic potential.
Figure 7: Stellar velocity dispersion versus age (top: radial velocity dispersion; bottom: vertical velocity dispersion). The blue line and band show the median and 16th to 84th percentile range from our simulations, computed as described in the main text. The orange line shows the power law relation from Sharma et al. (2021), rescaled and extrapolated to the younger ages in our simulation, while the green and red lines show the lower and upper limits, respectively, of the scaled power law relations found by Aumer et al. (2016) in their simulations. The $\beta$ values for each line are given in the legend.
4.2 Age-velocity dispersion relation
Stellar velocity dispersions are closely related to the distributions of stellar actions, and thus the age-velocity dispersion relation (AVR) in our simulations also offers the opportunity for a (slightly less direct) test of action evolution. We therefore compare the AVR in our simulations to two previous studies: the observational results of Sharma et al. (2021) and the NN-body simulations of Aumer et al. (2016). However, we caution that, unlike in our analysis of the action distribution from Garzon et al. (2024) in Section 4.1, this comparison can only be rough and qualitative, because the data sets available are much less well-matched to our simulations. Sharma et al.’s observational study contains no measurements for stars younger than 2 Gyr, and so to compare to our study over $\approx 0.5$ Gyr we must extrapolate their trends well beyond the range of the data. Similarly, Aumer et al. do not include star formation and stellar feedback, and include gas only in a subset of their simulations and only then with a fixed temperature of $10^{4}$ K rather than realistic heating and cooling; they rely on a subgrid model to attempt to capture the effects of the omitted physics. We should therefore not necessarily expect good agreement between their results and ours.
With these caveats understood, we start by computing the AVR in our simulation as follows: we first filter the stellar velocities using the same Butterworth filter we apply to actions (Section 3.2) to remove short-timescale oscillations associated with virial motion in bound clusters. Next, for each simulation snapshot, we compute the age $\tau$ of each star in that snapshot, and bin the stars by age in 1 Myr-wide bins. Within each bin, we compute the radial and vertical velocity dispersion, yielding values of $\sigma_{R,z}$ versus $\tau$ for that snapshot. Repeating this procedure for every snapshot yields multiple measurements of $\sigma_{R,z}(\tau)$, with the number of measurements decreasing with $\tau$ – for $\tau=1$ Myr we have 464 measurements (since each output snapshot includes stars that are 010-1 Myr old), while for $\tau=464$ Myr we have only a single measurement (since only the final output contains stars 463464463-464 Myr old). From this set of measurements, we compute the 16th, 50th and 84th percentiles of the resulting $\sigma_{R,z}$ distributions, which we plot as a function of $\tau$ in Figure 7.
For the purposes of comparing to the results from Aumer et al. (2016) and Sharma et al. (2021), we note that these authors fit the AVR to an empirical power law functional form
$\sigma_{R,z}(\tau)=C_{R,z}\left(\frac{\tau+0.1\,\mathrm{Gyr}}{10.1\,\mathrm{Gyr}}\right)^{\beta_{R,z}}$ (15)
where $C_{R,z}$ is a scaling constant and $\beta_{R,z}$ is the corresponding power-law exponent. The scaling constant is not particularly meaningful, since for Aumer et al. this is set primarily by their subgrid model (in which the initial stellar velocity dispersion is a free parameter), while for Sharma et al. our simulations are well outside the range of ages they have measured. For this reason, we rescale the values of $C_{R,z}$ found by these authors to match our results at $\tau=0$, allowing us to focus on the trend with time rather than the absolute values. We plot the resulting rescaled curves in Figure 7; for Aumer et al., we plot two curves, corresponding to the minimum and maximum values of $\beta$ they determine by varying the parameters of their subgrid models.
We find that the velocity dispersions of our simulation grow more slowly with age than in either comparison studies. For $\sigma_{z}$ the measured result from Sharma et al. is intermediate between the results of our simulations, which predict slower $\sigma_{z}$ growth, and the N-body simulations of Aumer et al., who predict faster-than-measured growth. The slower growth of the AVR in our simulations compared to the observational results from Sharma et al. (2021) may reflect the fact that their fits span $\approx 10$ Gyr, over which time the Milky Way was likely significantly more dynamically active – due to higher gas content and merger activity – than in the present-day state that we simulate. Our differences from Aumer et al. almost certainly arise due to the differences between our self-consistent treatment of gas and stellar feedback versus their subgrid model.
The most important point to take away from Figure 7, however, is that our simulations are quite dynamically quiet, with slow growth of the AVR. Significant changes appear to require Gyr timescales, here as well as in the previous studies. In Section 5, we contrast this slow, population-averaged evolution with the evolution of the individual stellar actions, which we show occurs on a much shorter timescale.
5 Results
For characterising the time-evolution of stellar actions, we first define some statistical measures of the same at the beginning of Section 5.1. We then study the properties of the time series describing individual stars rather than the full population. We discuss how the evolution varies with the environmental conditions under which stars are born in Section 5.2 and Section 5.3. In Section 5.4, we compare the action evolution for two different population of stars.
5.1 Change in actions over time
@@
where $i=(R,z,\phi)$ is the component, $t$ is the time of the snapshot, $\Delta t$ is the time interval between the two snapshots and $J_{i}(t)$ is the $i$ component of the star’s action action at time $t$ in the simulation. Similarly, we define the relative difference [^1] Astute readers might worry that computing relative difference in action could yield infinite values for stars with very small initial actions (e.g., those born on nearly circular orbits with $J_{R}\simeq 0$). While this is true in principle, our analysis avoids this problem by using median statistics, which are robust to such outliers and hence, we find that the relative changes remain well-behaved. as
$\delta J_{i}(t,\Delta t)=\frac{\Delta J_{i}(t,\Delta t)}{J_{i}(t)}.$ (17)
@@
We have $\approx 450$ time snapshots, and therefore $\approx 10^{5}$ snapshot pairs, and $\sim 10^{6}$ stars, so our total sample for analysis consists of $\sim 10^{11}$ action differences. While this gives us a very large sample size for statistical robustness, it is important to note that these measurements are not all independent as they include correlated measurements of the same stars at different times.
We are interested in how stellar actions change with time difference $\Delta t$, but also with stellar age. To facilitate this, we note that each star has a formation time $t_{\mathrm{form}}$, and for each action pair $(t,\Delta t)$ we define $t_{*}=t-t_{\mathrm{form}}$ as the age of the star at the earlier time – thus for example if we are computing differences in action between snapshots at 100 Myr and 250 Myr, and we are examining a star that formed 95 Myr into the simulation, then we have $t=100$ Myr, $\Delta t=150$ Myr, and $t_{*}=5$ Myr. Thus our action differences $\Delta J_{i}$ or $\delta J_{i}$ have both a time gap $\Delta t$ and an initial stellar age $t_{*}$ associated with them, and below we will bin our data by $\Delta t$ and $t_{*}$.
@@
Figure 8: Heatmap of the median of the absolute change in stellar actions (row-wise in order: $\Delta J_{R}$, $\Delta J_{z}$, and $\Delta J_{\phi}$) as a function of the time lag ($\Delta t$) and initial stellar age ($t_{*}$). The dotted lines indicate initial stellar ages of 5, 10, 50, 100, 200, 300 and 400 Myr, corresponding to the time slices plotted in Figure 9. The empty triangle in the top right of the plot corresponds to combinations of initial stellar age and time lag that are inaccessible due to the 464 Myr duration of the simulation that we analyse.
We now investigate the evolution of individual stellar actions over time as characterised by the absolute changes $\Delta J_{R}$, $\Delta J_{z}$ and $\Delta J_{\phi}$ (equation 16). These quantities are functions of both the stellar age at the earlier snapshot, $t_{*}$, and the time lag between snapshots, $\Delta t$. To analyse this, we place all stars in 2D bins of these quantities and compute the median [^2] Note that since we are taking the modulus of the difference in equation 16 and then taking median here, this is similar to taking a square root of the mean of the square of the difference in actions at the two times. in each bin; we use medians rather than means to ensure robustness against outliers. Before presenting the results, we note an important consequence of our definitions of $\Delta t$ and $t_{*}$. At small values of $t_{*}+\Delta t$, all stars formed during the simulation contribute to our analysis. However, only the stars formed near the start of the simulation contribute to the analysis at the limit $t_{*}+\Delta t<464$ Myr (which is the time for which we run the simulations, excluding the initial transient phase that we discard).
@@
We show the binned changes in absolute action in Figure 8; in this plot, initial stellar age $t_{*}$ appears on the $x$-axis and the time lag $\Delta t$ on the $y$-axis. The expected behaviour is immediately visible – small variations at short time intervals appear as the darkest colours in a horizontal band at the bottom of the plot. We see only a weak dependence on initial stellar age, as there are no very obvious vertical features in the heatmap.
@@
To take a closer look, we extract vertical slices from the heatmap at $t_{*}=5$, 10, 50, 100, 200, 300 and 400 Myr, as shown by the dashed lines in Figure 8. These slices, illustrated in Figure 9, show how the absolute change in actions depends on the time lag for stars of different initial ages. We see here that there is a weak dependence of $\Delta J$ on initial stellar age, but in a direction that is opposite what one might naively expect – we expect younger stars to exhibit larger changes in their actions than older stars because they are more likely to reside near their natal molecular clouds (Lada & Lada, 2003), where ongoing interactions with dense gas or other stars can lead to greater dynamical perturbations. Conversely, older stars have typically migrated away from their birth regions and are less influenced by such perturbations. Contrary to these expectations, we instead find that older stars undergo more rapid changes to their actions, as indicated by the larger slopes in Figure 9.
@@
Figure 9: Median of the absolute change in stellar actions (row-wise in order: $\Delta J_{R}$, $\Delta J_{z}$, and $\Delta J_{\phi}$) for stars of different ages across the simulation as a function of time lag. Different initial stellar age groups are shown in different colours as indicated in the legend.
@@
To understand why this occurs, we next examine relative rather than absolute changes in actions, focusing on $J_{R}$ and $J_{z}$ since $J_{\phi}$ is dominated by circular orbital motion and is less sensitive to dynamical perturbations. Figure 10 shows the median square relative changes of these two actions, $\delta J_{R}^{2}$ and $\delta J_{z}^{2}$, as a function of time lag for the same initial stellar ages as shown in Figure 9. In this plot, we see that the different initial stellar age bins collapse onto nearly a single line, suggesting a universal behaviour among stars of all ages.